Water (density = 1.00 ? 10 3 kg/m 3 ) flows at 10.0 m/s through a pipe with radius 0.030 m. The pipe goes up to the second floor of the building, 2.0 m higher, and thepressure remains unchanged. What is the radius of the pipe on the second floor?
Answer #1

By Bernoullis equation P/ρ + v2/2 + gh = constant for steady flow

therefore P1/ρ + v12/2 + gh1=P2/ρ + v22/2 + gh2

where 1 stands for bottom and 2 for top

given P1= P2

therefore v12-v22= 2gΔh

v22= 102- 2x9.8x2

v2= 7.8 m/s

Now by conservation of mass,ρAv = constant

A1v1=A2v2

πr12x 10 =πr22x 7.8

0.032x 10 = r22x 7.8

r2= 0.034 m

answered by: Adelle
Answer #2

ussing the kinematic equation,

v^2 = u^2 + 2gh ( where v= speed of water at second floor),( u =speed of lighht in first floor_)

v^2 = 10^2 - 2(9.8)(2)

v^2 = 100 - 39.2

v= 7.797 m/s apprx

using the equation of conmtinity

A1V1= A2V2

π( 0.030)^2 ( 10 ) =πr2( 7.797)

r = 0.034 m apprx---------radius at second loor

answered by: Charlie